A. Tell whether each given function has a solution on the indicated closed interval. Prove using the Intermediate value theorem.


1. f(x)=3x²+2x²;[-1,1]


2. f(x)= 2-x²/x²;[-3,-1]



B. Sketch the graph of the following functions and then find the absolute extreme values of each of the given interval.


1.f(x)=√x²-25;[5,10]


2. f(x)=-1/x²;[0.5,2]



C. A restaurant's profit function (in hundreds) for hamburgers is given by the function P such that P(x)=1.22x-x²/30,000-4,000,where 0≤x≤20,000.


1. How many hamburgers does the restaurant need to sell to yield the maximum profit?


2. What is the maximum profit from the sale of hamburger ?

1
Expert's answer
2022-03-19T02:43:02-0400

Solution (A)


Given that


"f(x)=3x^3+2x" with in interval "[-1,1]"


"f(-1)=3(-1)^3+2(-1)=-5<0"


"f(1)=3(1)^3+2(1)=5>0"


Since


"f(-1)<0" and "f(1)>0"


Therefore, according to IVT the function "f(x)=3x^3+2x" has a solution within the interval "[-1,1]"





The plot above shows that the function "f(x)=3x^3+2x" has a solution within the interval "[-1,1]", which is "x=0"



Solution (B)


for the function


"g(x)=\sqrt{x^2-25}" with in the interval "[5, 10]" , the plot is shown below





From the plot, the extreme values are


The maximum value is "8.66" at "x=10"


The minimum value is at "x=5"




Now For the function


"h(x)=-\frac{1}{x^2}" within the interval "[0.5,2]"


The plot is shown below





From the plot, the extreme values are


The maximum value is "-0.25" at "x=2"


The minimum value is "-4" at "x=0.5"


Solution (C)

"p(x)=1.22x-\frac{x^2}{30000}-4000" with in the interval "[0, 20000]" , the plot is showb below





From the plot,


The maximum profit is 7163


And the hamburger sold were 18300